OpenStudy (anonymous):

Of the athletes at My schoool, Greenville High School, 24 are on the soccer team and 34 are not. What is the probability that a randomly selected athlete is on the soccer team? Simplify your answer and write it as a fraction or whole number. @dmezzullo

OpenStudy (anonymous):

Lol. I live in a place called Greenville..Odd.

OpenStudy (meganfox):

Girl you remind me of me when I was young c:

OpenStudy (anonymous):

whats the total number of athlete?

OpenStudy (anonymous):

uhhh, ummm, cough, total is 24+34 = 58 P = 24/58 = 12 / 29